``find the value of m if 3 to the the power of 3 times 9 to the power of 3 times 27 to the power of 3 times ....19683 to the power of 3 =3 to the power of m.''.
please help

Answer by jorel1380(3719)

I believe your equation is:3^m=19683...so find m. If we take the natural log of both sides, we get:
ln 3^m=ln 19683
m ln 3=ln 19683
m = ln 19683/ln 3
m=9
Thus 3^9=19683. ☺☺☺☺